OpenStudy (anonymous):

Find the slope and the y intercept of the line whose equation is 2x+y=-4

OpenStudy (anonymous):

can you put the equation in this form y = mx + b just solve for y

OpenStudy (anonymous):

is -4 the slope?

OpenStudy (anonymous):

I am so confused

OpenStudy (nubeer):

no.. the "m" part in the equation would be slope.

OpenStudy (anonymous):

2x+y=-4 subtract 2 x from both sides 2x + y - 2x = -4 - 2x y = -2x -4 slope intercept form y = mx + b m is the slope do u kno the slope now?

OpenStudy (anonymous):

I SEEE!!! -2 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

THANK YOU!!!!!

OpenStudy (anonymous):

now plug in x = 0 to get the y-intercept

OpenStudy (anonymous):

wanna try it?

OpenStudy (anonymous):

ok.... so when that is plugged in the y intercept would be -4

OpenStudy (anonymous):

(0,-4)

OpenStudy (anonymous):

I see!! Thank you for your help

OpenStudy (anonymous):

no problem